Find in degrees the numeric value of the acute angle of rotation that eliminates the product term from the equation
7
x
2
+
24
x
y

34
x
+
24
y

185
=
0
1
Expert's answer
2020-06-03T19:23:39-0400

"7x^2+24xy-34x+24y-185=0"

The general equation of line 2 order:

"a_{11}x^2+2a_{12}xy+a_{22}y^2+2a_{1}x+2a_2y+a=0"

The acute angle of rotation eliminates the product term from the equation 

"\tan2\alpha=\frac{2a_{12}}{a_{11}-a_{22}}"

In our case

"a_{11}=7, 2a_{12}=24, a_{22}=0"

"\tan2\alpha=\frac{24}{7-0}=\frac{24}{7}=3.43\\
2\alpha=74^0\\
\alpha=37^0"